OpenStudy (anonymous):

if f(x)=1-x/1+x,x>0 then least value of f(f(x)) + f(f(1/x)) is what

OpenStudy (anonymous):

need some paper for this one

OpenStudy (anonymous):

is if its own inverse? looks like it

OpenStudy (anonymous):

so \[f(f(x))=x\]

OpenStudy (anonymous):

and likewise \[f(f(\frac{1}{x})=\frac{1}{x}\]

OpenStudy (anonymous):

so you are trying to minimize \[x+\frac{1}{x},x>0\]

OpenStudy (anonymous):

minimum is 2 if x = 1

OpenStudy (anonymous):

you need the method for this or is it clear?

OpenStudy (anonymous):

i need the method

OpenStudy (anonymous):

ok first is it clear that \[f(f(x))=x\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the fuction is its own inverse??

OpenStudy (anonymous):

then it should also be clear that \[f(f(\frac{1}{x}))=\frac{1}{x}\]

OpenStudy (anonymous):

well yes, if \[f(f(x))=x\] that means the function is its own inverses

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

understood thnx

OpenStudy (anonymous):

ok so our real job once the algebra is out of the way, is to find the minimum value of \[x+\frac{1}{x}\] for x > 0 there are two ways to do this, one is calculus and one is thinking. is this a calc course?

OpenStudy (anonymous):

We can use AM>=GM

OpenStudy (anonymous):

i guess so. if you use calc you take the derivative, find that it is \[1-\frac{1}{x^2}\] find the critical points where the derivative is 0 i.e. solve \[1-\frac{1}{x^2}=0\] \[\frac{1}{x^2}=1\] \[x=\pm1\]and since we know x> 0 we know x = 1, and we also know it is a minimum, and \[1+\frac{1}{1}=2\]

OpenStudy (anonymous):

think method is that this is symmetric in that if you replace x by 1/x you get the same thing back again

OpenStudy (anonymous):

meaning the minimum value has to occur when x and 1/x are equal, meaning they are both 1, and so you get 2

OpenStudy (anonymous):

and now that you mentioned it, using AM >= GM you see that \[\frac{1}{2}(x+\frac{1}{x})\geq \sqrt{x\times \frac{1}{x}}\] \[\frac{1}{2}(x+\frac{1}{x})\geq \sqrt{1}=1\] and so \[x+\frac{1}{x}\geq 2\]

OpenStudy (anonymous):

and since we can make it 2, by making x = 1 we know that this is a minimum i have to say i would not have thought of that one!!